If 15% of the 250 million adult Americans live in poverty, how many adult Americans live in poverty?
Answer by lenny460(1073)

If 15% of the 250 million adult Americans live in poverty, how many adult Americans live in poverty?

10% of 250 million = 25 million
5% = 25/2 = 12.5 million
Therefore 15% = 25 + 12.5 = 37.5 million

The Answer:
37.5 million